# EA4LLM: A Gradient-Free Approach to Large Language Model Optimization via Evolutionary Algorithms

### v1

@@ Abstract
To address these limitations, we propose a method for optimizing LLMs using evolutionary algorithms (EA4LLM) and, for the first time, successfully demonstrate its capability to train a 1-billion-parameter LLM from the pre-trained stage. We conduct extensive experiments and provide key insights into how evolutionary algorithms can effectively optimize neural networks.
@@
Figure 1: Loss Curve of EAGPT-1B Optimized from Pre-Trained Start Using Evolutionary Algorithms.
@@ 1 Introduction
Igel et al. (Igel, 2003) and Tim Salimans et al. (Salimans et al., 2017) explore using Evolution Strategies (ES) as an alternative to Markov Decision Process (MDP)-based methods for optimizing neural networks in reinforcement learning. In recent years, some methods have utilized EAs to optimize large models’ prompts (Sun et al., 2022a, b; Zhao et al., 2023; Guo et al., 2023) and small parameter spaces (Jin et al., 2024; Carmona et al., 2024; Huang et al., 2025), validating the ability of EAs to optimize large model parameters in smaller-scale settings. More recently, Xin Qiu et al. (Qiu et al., 2025) adapted the reward computation from GRPO  (Guo et al., 2025) to design the fitness evaluation and population update mechanism in their ES algorithm, successfully extending ES to full-parameter fine-tuning of large language models (LLMs) in the reinforcement learning stage for the first time. However, these works only apply ES within the reinforcement learning context, and their effectiveness heavily depends on the design of the RL reward function.
We further investigate the capability of Evolution Strategies for direct neural network optimization beyond reinforcement learning. We propose EA4LLM, a method that leverages evolutionary algorithms to optimize LLMs, and demonstrate for the first time that ES can effectively perform full-parameter optimization of LLMs starting from pretraining, as illustrated in 1. In our implementation, we introduce a simple yet effective approach that links the model’s output logits to the fitness computation in ES. This implies that virtually any neural network with a probabilistic training objective $P$ can, in principle, be optimized using ES. Different from prior ES in supervised learning that typically treats mini-batch loss or performance as a black-box objective, we explicitly aggregate token-level log-probabilities (log-softmax) into the ES fitness and carry out full-parameter optimization for large-scale LLMs in pretraining; combined with antithetic sampling, rank shaping, grouped/layered controls, and cross-process synchronized subsampled evaluation, this substantially improves the practicality and efficiency of zeroth-order optimization in this setting. Building upon standard ES (Salimans et al., 2017), we conduct extensive exploratory experiments and provide key insights—as well as identified shortcomings—regarding the direct application of ES to neural network optimization.
@@ 3.2 Evolution Strategies and Our Extensions
Antithetic Sampling.
To reduce variance, we employ antithetic (mirror) sampling. For base seeds $b_{k}$ we form pairs $\{+\epsilon_{b_{k}},-\epsilon_{b_{k}}\}$ and weight rewards with the corresponding sign before normalization:
$\tilde{r}_{j}=s_{j}\cdot r_{j},\qquad s_{j}\in\{+1,-1\},\quad j=1,\dots,N,$ (10)
then compute $w_{j}$ from $\{\tilde{r}_{j}\}$ (via z-score or rank shaping) and aggregate as above. This matches our implementation where signed rewards are normalized and used to weight the noise sum.
@@
Layered/Grouped Controls.
We partition parameters into semantic groups $g\in\mathcal{G}$, such as embeddings (emb), attention (attn), MLP (mlp), LayerNorm (ln), and output head (head). We introduce group-specific multipliers $\alpha_{g}$ and $\sigma_{g}$:
Perturbation: $\theta^{(g)}_{\text{pert}}=\theta^{(g)}+(\sigma\,\sigma_{g})\,\epsilon_{j}^{(g)},$ (13)
Update: $\theta^{(g)}\leftarrow\theta^{(g)}+\frac{\alpha\,\alpha_{g}}{s}\,\frac{1}{N}\sum_{j=1}^{N}w_{j}\,\epsilon_{j}^{(g)}.$
$s=\begin{cases}\sigma&\text{if using $1/\sigma$ scaling},\\ 1&\text{otherwise.}\end{cases}$ (14)
This mirrors our code: perturbations are scaled by group-specific $\sigma_{g}$, and updates use an effective step size $\alpha_{\mathrm{eff}}^{(g)}=(\alpha\,\alpha_{g})/s$.
Schedulers.
We further employ step-decay schedulers for $\alpha$ and $\sigma$ to control optimization dynamics. Let iteration index be $t$. With decay step sizes $S_{\alpha},S_{\sigma}$ and decay factors $\gamma_{\alpha},\gamma_{\sigma}$:
$\alpha_{t}=\alpha_{0}\,\gamma_{\alpha}^{\left\lfloor t/S_{\alpha}\right\rfloor},\qquad\sigma_{t}=\sigma_{0}\,\gamma_{\sigma}^{\left\lfloor t/S_{\sigma}\right\rfloor}.$ (15)
@@ 4.1 Experimental Setups
  • Model: Qwen3 ~1B configuration (see Appendix A.4).
@@
Switch-Based Experimental Variants.
we systematically compare the following configurations:
  • Baseline: vanilla ES without antithetic, $1/\sigma$ scaling, rank shaping, or grouped multipliers.
  • Antithetic: enable mirror sampling ($s_{j}\in\{\pm 1\}$) to reduce estimator variance.
  • Antithetic + $1/\sigma$: apply NES-style scaling, i.e., $\theta\leftarrow\theta+(\alpha/\sigma)\hat{g}$.
  • Antithetic + $1/\sigma$ + Rank Shaping: replace z-score with $w_{j}^{\text{rank}}\in[-1,1]$ to be robust to outliers.
  • Layered $\alpha_{g}$: set per-group multipliers (e.g., embeddings <1<1, attention/MLP $\approx 1$, LN/head <1<1) to modulate update magnitudes across modules.
  • Layered $(\alpha_{g},\sigma_{g})$: additionally apply per-group $\sigma_{g}$ during perturbation to control noise scale by layer type.
  • Schedulers: introduce $\alpha_{t}$ and $\sigma_{t}$ step decays with $(\gamma_{\alpha},S_{\alpha})$ and $(\gamma_{\sigma},S_{\sigma})$ to gradually anneal step size and noise.
Across all variants, we aggregate per-seed rewards over the evaluation texts, normalize them (z-score or rank), form $\hat{g}$ by summing seeded noises weighted by $w_{j}$, and update all parameters—then synchronize weights across evaluation threads. These extensions preserve the ES estimator while improving stability, variance reduction, and layer-wise controllability for LLM pretraining from random initialization.
@@ 5 Conclusion
Our work introduces a practical, architecture-agnostic zeroth-order training path for large language models that is consistent with maximum likelihood. By explicitly aggregating token-level log-softmax into an ES fitness, we optimize LLMs from the pretraining phase without backpropagation. This design reduces reliance on gradient storage and differentiability, opening opportunities for models and systems that include non-differentiable components. On the engineering side, antithetic sampling, z-score/rank shaping, grouped/layered multipliers for $\alpha$ and $\sigma$, and cross-process synchronized subsampled evaluation together deliver stable updates and tractable cost at scale. Conceptually, our approach bridges probabilistic modeling objectives with ES estimators, yielding an unbiased mini-batch estimator under uniform subsampling and a population-based update that is robust to outliers and scale. Practically, it lowers the barrier for low-resource teams to perform meaningful LLM training, and provides a foundation for hybrid pipelines (e.g., alternating ES and gradient steps) and black-box optimization scenarios.
The Limitations of Our Work
While promising, our method has limitations. First, ES estimators can exhibit higher variance than first-order gradients; performance is sensitive to population size, noise scale $\sigma$, learning rate $\alpha$, and the choice of normalization (z-score vs. rank shaping), requiring careful scheduling and tuning. Second, sample efficiency may be lower than gradient-based training: each iteration needs multiple forward evaluations, and although subsampling reduces cost, it increases stochasticity even if unbiased. Third, using token-level log-likelihood as fitness may not align with certain downstream metrics (e.g., factuality, instruction-following) unless weighting or task-specific shaping is introduced. Fourth, grouped/layered multipliers and schedulers are currently heuristic and architecture-dependent; generalization across model families remains to be validated. Fifth, theoretical guarantees in non-convex settings are limited compared to well-studied gradient methods. Finally, scalability to extremely large models depends on efficient parameter perturbation, noise broadcasting, and I/O; implementation complexity and integration with standard training stacks can be nontrivial.
Future Directions
We envision several avenues to strengthen this line of work: adaptive or learned schedules for $\alpha$ and $\sigma$, variance reduction beyond antithetic sampling (e.g., control variates), task-aware fitness shaping and masking, hybrid ES/gradient training schemes, improved grouping strategies informed by layer sensitivity, and more comprehensive theory and benchmarks at LLM scale.
@@ A.1 Differences from ES in Supervised Learning
Salimans et al. (2017) studied, on small supervised datasets (e.g., MNIST), the correlation between ES gradient estimates $g_{t}^{ES}$ obtained via random perturbations and the true gradients $g_{t}^{True}$ from backpropagation, using the mini-batch performance/loss as a black-box objective $f(\theta+\sigma\epsilon)$. Although that objective is computed from logits, it does not explicitly structure token-level log-probabilities as a reusable ES fitness, nor does it provide a systematic, full-parameter ES optimization for large-scale LLMs in pretraining.
Our approach explicitly rewrites the language modeling maximum-likelihood objective as ES fitness. Let the model outputs (logits) be $z_{t}$. We define
@@
where $m_{t}$ can be used for masking or weighting, and $F(\theta)$ aligns with the standard log-likelihood objective in language modeling. We use $F$ as the ES fitness during pretraining and combine antithetic sampling, rank shaping, grouped/layered noise and step-size controls, and cross-process synchronized subsampled evaluation, making zeroth-order optimization feasible and efficient at LLM scale.
@@ A.2 Differences from ES in the RL Stage
This line of work centers on GRPO-style rewards and is tailored to RL fine-tuning, where the objective is a task-specific reward function. Its performance heavily depends on reward design and the sampling environment, and it does not directly target the pretraining objective of language modeling.
We apply ES to full-parameter optimization starting from pretraining, with fitness defined as the aggregation of token-level log-probabilities consistent with maximum likelihood, independent of external environments or rewards. Grouped and scheduled controls further enable stable updates at large model scales.
@@ A.3 Differences from MeZO and Other Zeroth-Order Optimizers
MeZO is a memory-efficient zeroth-order optimizer that typically uses two function evaluations $f(\theta\pm\delta u)$ to approximate directional derivatives and then updates parameters in a first-order style, with the objective treated as a black-box mini-batch loss. In contrast, we:
  • employ population-based ES with antithetic sampling, aggregating multiple perturbation directions per iteration;
  • explicitly aggregate token-level log-softmax as ES fitness, and use z-score/rank shaping to reduce scale sensitivity and outlier effects;
  • introduce grouped/layered noise and step-size multipliers (with scheduling of $\alpha$ and $\sigma$) for controllability and stability;
  • use cross-process synchronized subsampled evaluation to reduce cost while preserving unbiasedness $\mathbb{E}_{\mathcal{D}_{t}}[F_{M}]=F$.
Therefore, our method does not merely treat the mini-batch loss as a black-box; it integrates the structured probabilistic objective of language modeling directly into ES fitness, together with stabilization and engineering strategies for large-scale pretraining.
@@ A.4 Model Configuration Details
We instantiate a Qwen3-style causal language model using the Transformers implementation (Qwen3Config) with the following settings, chosen to approximate a $\sim$1B-parameter configuration while tying input and output embeddings:
  • Vocabulary size: 151,936151{,}936.
  • Hidden size: 1,5361{,}536.
  • Intermediate size: 4,0964{,}096.
  • Number of layers: $24$ transformer blocks.
  • Attention heads: $12$; KV heads: $4$ (GQA).
  • Head dimension: $128$.
  • Max position embeddings: 32,76832{,}768.
  • Attention bias: false.
  • Tie word embeddings: true (output head tied to input embeddings).
  • Use cache: false (disabled during ES evaluation).
  • Attention dropout: $0.0$.
This configuration aligns with the training scripts used in our experiments and balances model capacity, context length, and ES evaluation stability.

### v2

@@ Abstract
To address these limitations, we propose EA4LLM, an evolutionary algorithm for optimizing LLMs, and, for the first time, empirically verify full-parameter optimization from the pretraining stage across model sizes ranging from 0.5B to 32B. We conduct extensive experiments and provide key insights into how evolutionary algorithms can effectively optimize neural networks.
@@
Figure 1: Loss Curve of EAGPT Optimized from Pretraining Using Evolutionary Algorithms.
@@ 1 Introduction
Igel et al. (Igel, 2003) and Tim Salimans et al. (Salimans et al., 2017) explore using Evolution Strategies (ES) as an alternative to Markov Decision Process (MDP)-based methods for optimizing neural networks in reinforcement learning. In recent years, some methods have utilized EAs to optimize large models’ prompts (Sun et al., 2022a, b; Zhao et al., 2023; Guo et al., 2023) and small parameter spaces (Jin et al., 2024; Carmona et al., 2024; Huang et al., 2025), validating the ability of EAs to optimize large model parameters in smaller-scale settings. More recently, Xin Qiu et al. (Qiu et al., 2025) adapted the reward computation from GRPO  (Guo et al., 2025) to design the fitness evaluation and population update mechanism in their ES algorithm, successfully extending ES to full-parameter fine-tuning of LLMs in the reinforcement learning stage for the first time. However, these works only apply ES within the reinforcement learning context, and their effectiveness heavily depends on the design of the RL reward function.
We further investigate the capability of Evolution Strategies for direct neural network optimization beyond reinforcement learning. We propose EA4LLM, a method that leverages evolutionary algorithms to optimize LLMs, and demonstrate for the first time that ES can effectively perform full-parameter optimization of LLMs starting from pretraining—validated across model sizes from 0.5B to 32B—as illustrated in 1. In our implementation, we introduce a simple yet effective approach that links the model’s output logits to the fitness computation in ES. This implies that virtually any neural network with a probabilistic training objective $P$ can, in principle, be optimized using ES. Different from prior ES in supervised learning that typically treats mini-batch loss or performance as a black-box objective, we explicitly aggregate token-level log-probabilities (log-softmax) into the ES fitness and carry out full-parameter optimization for LLMs in pretraining; combined with antithetic sampling, rank shaping, grouped/layered controls, and cross-process synchronized subsampled evaluation, this substantially improves the practicality and efficiency of zeroth-order optimization in this setting. Building upon standard ES (Salimans et al., 2017), we conduct extensive exploratory experiments and provide key insights—as well as identified shortcomings—regarding the direct application of ES to neural network optimization.
@@ 3.2 Evolution Strategies and Our Extensions
Algorithm 1 EA4LLM: Zeroth-Order Pretraining via ES
Input: Pretrained LLM parameters $\theta_{0}$; token-level fitness $R(x;\theta)$ via log-softmax; iterations $T$; population size NN; noise scale $\sigma$; learning rate $\alpha$; subset size $M$; normalization (z-score or rank shaping).
for $t=1$ to $T$ do
  Sample a uniform subset $\mathcal{D}_{t}\subset\mathcal{D}$ with $|\mathcal{D}_{t}|=M$ and broadcast indices across processes.
  Draw $\{\epsilon_{j}\}_{j=1}^{N}$ with $\epsilon_{j}\sim\mathcal{N}(0,I)$.
  for $j=1$ to NN do
   Form temporary parameters: $\theta_{\text{tmp}}=\theta+\sigma\,\epsilon_{j}$.
   Evaluate reward on the subset: $r_{j}=\frac{1}{M}\sum_{x\in\mathcal{D}_{t}}R\big(x;\theta_{\text{tmp}}\big)$.
  end for
  Compute weights $w_{j}$ from $\{r_{j}\}$ by z-score normalization or rank shaping.
  Aggregate noise: $\hat{g}=\frac{1}{N}\sum_{j=1}^{N}w_{j}\,\epsilon_{j}$.
  Update parameters: $\theta\leftarrow\theta+\alpha\,\hat{g}$.
end for
Return: $\theta_{T}$.
The pseudocode of our final algorithm is shown in Algorithm 1.
@@ 4.1 Experimental Setups
  • Models: Qwen3 configurations spanning 0.5B–32B (see Appendix A.4).
@@
4.2 Experimental Results
Cross-Scale Training Validation (0.5B–32B).
We pretrain Qwen3-style models from 0.5B to 32B with full-parameter updates and consistently observe stable loss reduction; the trend matches the loss curve shown earlier in Figure 1, indicating ES effectively optimizes across model scales.
Memory Footprint Comparison: Evolutionary vs Gradient Optimization.
We compare the minimal GPU memory usage when training 0.5B–32B models using ES versus gradient-based pipelines. By avoiding backpropagation and the storage of gradients/optimizer states, ES exhibits lower memory usage across all scales; the relative advantage increases with model size. Moreover, as model parameters increase, the memory advantage becomes more pronounced; for models of 4B and above, ES-based pretraining reduces GPU memory usage by more than $4\times$ compared to gradient-based methods. See Figure 2.
Figure 2: Minimal GPU memory usage across model sizes (0.5B–32B), comparing ES with gradient-based optimization. ES consistently uses less memory, with a larger margin as the model size increases.
Ablations: Evaluation Budget and Population Size.
On the 1.7B model, we study the effect of evaluation samples per iteration (1000/3000/5000/7000) and population size (20/30/40). Results match expectations: larger evaluation budgets and larger populations provide more accurate optimization signals, leading to more pronounced loss reduction. Considering time and performance constraints, in other experiments we primarily adopt evaluation budget 5000 and population size 30. See Figure 3 and Figure 4.
Figure 3: Different evaluation budgets (1.7B). Larger budgets yield more accurate optimization signals and stronger loss reduction.
Figure 4: Different population sizes (eval=5000). Larger populations yield more accurate optimization signals and stronger loss reduction.
@@ 5 Conclusion
Our approach provides a practical, architecture-agnostic zeroth-order pretraining path for large language models by tying token-level log-softmax to an ES fitness (Section 3). We employ subsampled evaluation to reduce per-iteration cost while maintaining an unbiased estimator under uniform sampling, and aggregate perturbations via weights from either z-score or rank shaping. Experimentally (Section 4), ES delivers stable loss reduction from 0.5B to 32B with full-parameter updates, and achieves substantial memory advantages over gradient-based training: for models of 4B and above, ES-based pretraining reduces GPU memory usage by more than $4\times$. Ablations on a 1.7B model confirm expected trends: larger evaluation budgets and populations provide more accurate optimization signals and stronger loss reduction; in practice, we primarily adopt an evaluation budget of 5000 and a population of 30.
Limitations and Future Work
While promising, zeroth-order ES pretraining has limitations. Variance can be higher than first-order gradients and performance is sensitive to population size $N$, noise scale $\sigma$, learning rate $\alpha$, and normalization choice; subsampling, though unbiased, introduces additional stochasticity. Sample efficiency may trail gradient-based methods because each iteration requires multiple forward evaluations. Scaling to extremely large models depends on efficient parameter perturbation and IO. In our experiments, certain variance-reduction and control mechanisms (e.g., antithetic sampling, NES-style $1/\sigma$ scaling, grouped multipliers) were disabled; exploring these and hybrid pipelines (alternating ES and gradient steps) is a natural direction. Future work includes adaptive schedules for $\alpha$ and $\sigma$, improved variance reduction, and task-aware fitness shaping beyond next-token log-likelihood to better align with downstream objectives.
@@ A.1 Differences from ES in Supervised Learning
Salimans et al. (2017) studied, on small supervised datasets (e.g., MNIST), the correlation between ES gradient estimates $g_{t}^{ES}$ obtained via random perturbations and the true gradients $g_{t}^{True}$ from backpropagation, using the mini-batch performance/loss as a black-box objective $f(\theta+\sigma\epsilon)$. While that objective is computed from logits, it treats the loss as an opaque function and does not offer a full-parameter ES optimization of large-scale LLMs for pretraining.
Our approach explicitly reformulates the maximum-likelihood language modeling objective as ES fitness. Let the model outputs (logits) be $z_{t}$. We define
@@
where $m_{t}$ can be used for masking or weighting, and $F(\theta)$ aligns with the standard log-likelihood objective in language modeling. We use a simple population ES with i.i.d. Gaussian perturbations $\epsilon_{j}\sim\mathcal{N}(0,I)$; for each iteration we evaluate $F(\theta+\sigma\epsilon_{j})$, apply either z-score normalization or rank-based shaping to the rewards, estimate the ES direction, and update parameters with the core ES step $\theta\leftarrow\theta+\alpha\,\hat{g}$. This design removes antithetic sampling, grouped/layered operators, and step-size multipliers, matching the actual algorithm we employ for large-scale pretraining while remaining effective empirically.
@@ A.2 Differences from ES in the RL Stage
RL-stage ES and GRPO-style methods focus on task-specific rewards for fine-tuning and depend on reward design and sampling environments. In contrast, we apply ES from pretraining to full-parameter optimization with fitness defined by the aggregation of token-level log-probabilities consistent with maximum likelihood, independent of external environments or rewards. This directly targets the pretraining objective and avoids reliance on RL-specific reward shaping.
@@ A.3 Differences from MeZO and Other Zeroth-Order Optimizers
MeZO is a memory-efficient zeroth-order optimizer that typically uses two function evaluations $f(\theta\pm\delta u)$ to approximate directional derivatives and then updates parameters in a first-order style, with the objective treated as a black-box mini-batch loss. In contrast, our method:
  • uses population-based ES with i.i.d. Gaussian perturbations, without antithetic pairs or grouped/layered operators;
  • integrates the structured token-level log-softmax objective into ES fitness, rather than treating the loss as opaque;
  • stabilizes updates via z-score or rank-based reward normalization to reduce scale sensitivity and outlier effects;
[…]
Therefore, our method directly optimizes the structured probabilistic objective of language modeling with a simple ES update and practical normalization strategies, making zeroth-order pretraining viable at LLM scale.
@@ A.4 Model Configuration Details
We instantiate a series of Qwen3-style causal language models using the Transformers implementation (Qwen3Config) spanning 0.5B–32B parameters in our experiments. The key configuration settings for each size are summarized below.
Size Layers Hidden Heads KV Heads Head Dim Intermediate Max Pos Tie Emb Use Cache Vocab
0.5B 24 1024 8 4 128 4096 32768 true false 151936
1.7B 28 2048 16 8 128 6144 32768 true true 151936
4B 36 4096 32 8 128 9728 32768 true true 151936
8B 36 4096 32 8 128 12288 32768 false true 151936
14B 40 5120 40 8 128 17408 32768 false true 151936
32B 64 8192 64 8 128 25600 40960 false true 151936
Table 1: Qwen3-style model configurations used for ES pretraining experiments across 0.5B–32B.